OpenStudy (anonymous):

If a^x = b^y = c^z and b^2 = ac, then find 1/x + 1/z.

OpenStudy (anonymous):

a^x = c^z x log(a) = z log(c) x log(a) + z log(a) = z log(c) + z log(a) (x+z) log(a) = z ( log (a) + log(c) ) (x+z) log(a) = z log (ac) (x+z) log(a) = z log (b^2) as b^2=ac (x+z) log(a) = 2z log (b) as a^x = b^y, x log(a) = y log(b) log (b) = (x/y) * log(a) hence, (x+z) log(a) = 2z*(x/y) * log(a) (x+z) = 2xz/y (x+z)/xz = 2/y 1/x + 1/z = 2/y

OpenStudy (anonymous):

Does this have to involve logarithms? Because I'm just in ninth grade.

OpenStudy (anonymous):

Another method is Let ( a^x ) = ( b^y ) = ( c^z ) = k. hence, a = k^(1/x), b = k^(1/y), c = k^(1/z). as b² = ac. [ k^(1/y) ]² = [ k^(1/x) ]·[ k^(1/z) ] [ k^(2/y) = k^[ (1/x) + (1/z) ] 2/y = (1/x) + (1/z) 2/y = ( z+x) / (xz) ( z+x) / (xz) = 2/y 1/x + 1/z = 2/y what about this ?

OpenStudy (anonymous):

Thanks!!